OpenStudy (anonymous):

If f(x)=10^x show that f(x+h)-f(x) --------- = 10^x((10^h -1)/h) h

OpenStudy (anonymous):

did you try

OpenStudy (anonymous):

yes idk where to start

OpenStudy (anonymous):

f(x + h) = ?

OpenStudy (anonymous):

find it,

OpenStudy (anonymous):

and try to form the given expression

OpenStudy (anonymous):

f(x) = 10^x => f(x+h) = ?

OpenStudy (anonymous):

10^x+h